OpenStudy (anonymous):

why are these 2 limit equal?

OpenStudy (anonymous):

Which?

OpenStudy (anonymous):

@PHP

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} n^2\log(n*\arctan(\frac{ 1 }{ n }))\] and \[\lim_{x \rightarrow 0}\frac{ \log(\frac{ \arctan(x) }{ x }) }{ x^2 }\]

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):